The perimeter of a rectangular sheet is 100 cm. If the length is 35 cm, find its breadth.
Also find the area.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Perimeter of the a rectangular sheet = 100 cm
Length of the rectangular sheet = 35 cm
We know that,
Perimeter of the rectangle = 2 (Length + Breadth)
100 = 2 (35 + Breadth)
(100/2) = 35 + Breadth
50–35 = Breadth
Breadth = 15 cm
Then,
Area of the rectangle = Length × Breadth
= 35 × 15
= 525 cm^2
∴Area of the rectangular sheet is 525 cm^2

Thanks